OpenStudy (anonymous):

I need help in working out: 4 - 3 Square root 3 / 5 - 3 Square root 2

OpenStudy (anonymous):

\[ 4 - 3 \sqrt{2} OVER 5 - 3 \sqrt{2}\]

OpenStudy (anonymous):

use the calculator. 4-(18 sqrt(2))/5 =-1.091168824543142175686079407154913082850818751357013063436...

OpenStudy (anonymous):

I forgot to mention, I have to rationalise the denominators!

OpenStudy (anonymous):

you have to rationalise the denominator as follows 4 - 3 root2 5 + 3root2 4(5) + 4*3root2 - 5*3root2 - (3root2)^2 ---------- x ----------- = ----------------------------------- 5 - 3root2 5 + 3root2 (5)^2 - (3root2)^2

OpenStudy (anonymous):

this further gives 20 + 12root2 - 15root2 - 18 2 - 3root2 = --------------------------- = ----------- 25 - 18 7

OpenStudy (anonymous):

if u want u can write it as 2 3 --- - ---- root2 7 7